Assume that the variable X is distributed as a Poisson with λ = 5.7. Calculate the following probabilities.

a. Pr( X = 1 )

b. Pr( X < 1 )

c. Pr( X > 1 )

d. Pr( X ≤ 1 )
Answer #1

Solution :

Given that ,

mean = \lambda = 5.7

Using poisson probability formula,

P(X = x) = (e-\lambda * \lambdax ) / x!

a)

P(X = 1) = (e-4.3 * 5.71) / 1! = 0.0191

Probability = 0.0191

b)

P(X < 1) = P(X = 0) =  (e-4.3 * 5.70) / 0! = 0.0033

Probability = 0.0033

c)

P(X > 1) = 1 - P(X \leq 1) = 1 - [P(X = 0) + P(X = 1) ]

= 1 - 0.091 - 0.0033

= 0.9776

Probability = 0.9776

d )

P(X \leq 1) = P(X = 0) + P(X = 1) ] = 0.091 + 0.0033 = 0.0224